OpenStudy (anonymous):

How do I solve 2-√10/2 to get -.58

OpenStudy (astrophysics):

\[2- \frac{ \sqrt{10} }{ 2 }\] like this?

OpenStudy (astrophysics):

Or \[2- \sqrt{\frac{ 10 }{ 2 }}\]

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

Are you sure? I'm thinking \[\frac{ 2-\sqrt{10} }{ 2}\] (parentheses will make your life so much easier)

OpenStudy (astrophysics):

Ah right, it's always hard to tell what the expression is so please @cdelarosa1031 put parentheses on your questions or upload an image. So is it what Peachpi said or one of the ones I mentioned above?

OpenStudy (anonymous):

It's not giving me parentheses. It's just like that and want to know how they got -.58

OpenStudy (anonymous):

Astrophysics, the first one

OpenStudy (astrophysics):

No...we mean put parentheses in your problem so we know what the question is, is the whole thing being divided by 2 (2-sqrt(10))/2 or just the sqrt(10)

OpenStudy (astrophysics):

Ok cool

OpenStudy (astrophysics):

\[2- \frac{ \sqrt{10} }{ 2 } = \frac{ 2 }{ 1 } - \frac{ \sqrt{10} }{ 2 }\] we can deal with the sqrt(10) after I guess, so do you know how to subtract fractions?

OpenStudy (anonymous):

I haven't done that in a while :/

OpenStudy (usukidoll):

@peachpi 's version takes me to -.58

OpenStudy (astrophysics):

Yeah I think it is @peachpi version because 1 gives u 0.42

OpenStudy (anonymous):

How are you getting the decimal answers?

OpenStudy (usukidoll):

\[\frac{ 2-\sqrt{10} }{ 2} \rightarrow \frac{2-3.16}{2} \rightarrow \frac{-1.16}{2} \rightarrow -0.58\]

OpenStudy (anonymous):

I was skipping a step :/

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (astrophysics):

Please make your question clear next time to avoid confusion @cdelarosa1031

OpenStudy (anonymous):

How about this problem:

OpenStudy (anonymous):

1+3/y/5/y

OpenStudy (anonymous):

Not sure how to upload image to show proper way on paper